a number 1 to 9 digits it can be eg 1+2-34etc and the numbers should not be repeated but the answer should be =100


Answer by Edwin McCravy(20055)

The first 7 counting numbers 1,2,3,4,5,6,7 
average to their middle number 4, so they add to 7×4 or 28.
Then the next two, 8×9 = 72 and 28+72=100.  So:

1+2+3+4+5+6+7+8×9 = 100

           28+72  = 100

Edwin